# Single-site dissipation stabilizes a superconducting nonequilibrium steady state in a strongly correlated lattice

### v1

@@ Abstract
Can superconducting order be made an attractor of open-system dynamics? We answer this question in the affirmative by proposing a minimal dissipation-engineering protocol for the particle–hole symmetric Hubbard chain. We apply a rotated quantum jump, namely a locally rotated $\eta$-pair lowering operator, on as little as a single lattice site and show that the resulting Lindblad evolution pumps the system from the vacuum into a nonequilibrium steady state (NESS) with $\eta$-pair off-diagonal long-range order (ODLRO). In contrast to schemes that rely on extensive dissipation or cooling into low-temperature phases, here a strictly local dissipative seed is sufficient to induce macroscopic pairing coherence across the entire interacting lattice. We elucidate the underlying mechanism from three complementary perspectives: the local dark-state structure of the rotated jump, strong-coupling protection of the holon–doublon manifold, and a Liouvillian symmetry/invariant-subspace picture that exposes the attractive fixed point and its spectral gap. We further show that the $\eta$-paired NESS is robust against a broad class of static disorder, which mainly modifies transients and relaxation rates, while identifying specific perturbations that dephase the pseudospin coherence and suppress long-range order. Our results establish a disorder-tolerant route to stabilizing superconducting order as a nonequilibrium attractor via minimal, local quantum-jump engineering.
@@ I Introduction
The paradigm of quantum state engineering has shifted fundamentally in recent years: rather than treating dissipation solely as a mechanism of decoherence, it is increasingly viewed as a resource for autonomous state preparation and stabilization[20, 16, 6, 11, 24, 19, 29, 23, 10]. By engineering the coupling between a system and its environment, one can design Liouvillian dynamics whose attractor (dark state) is a desired quantum many-body state[26, 27, 22, 31]. This approach has successfully demonstrated the stabilization of entangled states[18, 3, 25], topological phases[13, 2, 17, 15], and superfluid correlations[12, 14, 9] in various atomic, molecular, and optical (AMO) and solid-state platforms.
However, a major open challenge lies in extending these protocols to strongly correlated lattice systems where competing orders and constraints make the phase diagram fragile. Most existing schemes for stabilizing long-range order rely on spatially extensive dissipation—engineering reservoirs that couple to every site or bond of the lattice[12, 29]. While theoretically powerful, such bulk engineering is experimentally demanding and often incompatible with the natural local addressing capabilities of quantum devices. This raises a fundamental question of control and scalability: Is it possible to stabilize a macroscopic, strongly correlated quantum phase using only a minimal, strictly local dissipative seed?
In this work, we answer this question in the affirmative by demonstrating the dissipative stabilization of off-diagonal long-range order (ODLRO) in the Hubbard model using a single local dissipator. We focus on the paradigmatic setting of $\eta$-pairing superconductivity[30], an exact eigenstate phenomenon in the Hubbard model that implies distance-independent pairing correlations protected by an SU(2) pseudospin symmetry. While $\eta$-pairing states are typically high-energy excited states in equilibrium and are thus thermodynamically unstable[28, 21], they represent ideal candidates for nonequilibrium steady states (NESS) if the appropriate dynamical symmetries can be enforced[5, 8].
We introduce a protocol based on a rotated quantum jump operator acting on as few as a single lattice site. Our results reveal a striking local-to-global mechanism: this local dissipative seed effectively pumps the entire interacting many-body system into a specific holon–doublon manifold, triggering a self-organization process that establishes global superconducting coherence. Crucially, this mechanism bypasses the need for bulk reservoir engineering. Instead, it exploits the interplay between the local dark-state structure enforced by the jump operator and the coherent Hamiltonian dynamics, which propagates the locally injected pseudospin coherence across the lattice.
@@
The remainder of this paper is organized as follows. In Sec. II, we introduce the heuristic principle of rotated dissipation using a single-qubit toy model. Section III formalizes this intuition in the many-body Hubbard lattice, providing a rigorous analysis of the symmetry-protected subspaces and the Liouvillian spectrum that enables the local-to-global mechanism. In Sec. IV, we present numerical evidence for the dynamical emergence of $\eta$-pairing ODLRO, tracking the real-time evolution of pairing correlators. Section V presents a comprehensive study of robustness, classifying disorder into benign renormalization effects versus destructive symmetry-breaking perturbations. Finally, we summarize our findings and discuss outlooks for experimental realization in Sec. VI.
Figure 1: Schematic of the nonequilibrium phase-locking protocol that stabilizes an $\eta$-paired superconducting NESS. Starting from a strongly correlated Hubbard system (with optional disorder in the interaction $U_{i}$, onsite potential $\mu_{i}$, and/or bond hopping $t_{ij}$), we apply local rotated dissipation on selected sites with jump operators $L_{j}=\eta_{j}^{-}e^{i\frac{\pi}{2}\eta_{j}^{x}}$. Unlike the unrotated channel $\tilde{L}_{j}=\eta_{j}^{-}$, which simply annihilates pairs and leaves the vacuum as the local dark state, the rotation tilts the dissipative measurement axis so that the locally protected state is the +y+y $\eta$-pseudospin eigenstate, $|\eta_{j}^{y}=+1/2\rangle\propto e^{-i\frac{\pi}{2}\eta_{j}^{x}}|0\rangle_{j}$, i.e., a coherent holon–doublon superposition. Each quantum jump then acts like a small compass correction: it damps misaligned components while repumping transverse $\eta$ coherence. In the presence of strong correlations, this local axis selection propagates through the many-body dynamics, effectively phase-locking the $\eta$ pseudospins across the lattice. At long times the system relaxes to a unique NESS with $\eta$ pseudospins aligned along the +y+y direction, yielding a finite $\eta$-pair amplitude and ODLRO.
@@ II Single-qubit heuristic: rotated dissipation pumps coherence from the vacuum
and set $H=0$ throughout this section. The goal of this toy example is to make one mechanism completely transparent, in a form that will reappear in the Hubbard chain: a jump operator does not merely relax the system, it selects an attractive fixed point (a dark state or dark manifold). Consequently, by rotating the jump one rotates the dissipatively selected axis and therefore rotates the steady state itself. In this way, dissipation can act as a simple coherence pump rather than a destroyer of phase information.
@@
Equivalently, the steady-state Bloch vector points along $\hat{y}$, with $\langle s^{x}\rangle=0$ and $\langle s^{y}\rangle=1/2$. Under the mapping (3), Eq. (6) is precisely the single-site analogue of $\langle\eta_{i}^{+}\rangle\neq 0$: the rotated dissipation does not “create coherence out of nothing”; rather, it reshapes the dark-state structure so that the unique attractor contains the desired coherence between $|0\rangle$ and $|\uparrow\downarrow\rangle$.
This local picture foreshadows the many-body mechanism exploited below. In the Hubbard chain, implementing the analogous rotated jump in the $\eta$-spin sector selects locally coherent holon–doublon superpositions as building blocks. The remaining task is to understand how interactions and hopping propagate and phase-lock this locally seeded $\eta$-pseudospin orientation across the lattice, resulting in an $\eta$-paired nonequilibrium steady state with off-diagonal long-range order.
@@ III.4 Leading projected Liouvillian, NESS, and the dissipative gap
From the perspective of potential objections, three clarifications are useful. (i) Effective gap and timescales: Eq. (26) should be read as a property of the projected effective Liouvillian, namely that the leading-order reduced generator possesses a finite spectral separation scale set by the local pumping strength. In the full microscopic model, additional processes (e.g., hopping-assisted excursions and local sector mixing) renormalize relaxation pathways and can obscure a clean single-exponential rate extraction; rather than over-interpreting a fitted microscopic gap, we use $\Delta\sim\gamma/2$ as an organizing scale that captures why convergence remains rapid and why deviations are suppressed in the strong-coupling/controlled window where the effective elimination applies. (ii) Finite-size scaling: to separate genuine long-range order from finite-size saturation, we track the $\eta$-pair structure factor $S_{\eta}(\tau)$ as well as the global amplitude $|\Phi(\tau)|$ for increasing system sizes (Fig. 3); the weak size dependence of the steady-state plateaus supports an $\mathcal{O}(1)$ order parameter in the thermodynamic limit. (iii) Uniqueness within the studied sector: starting from the vacuum (holon–doublon) sector, the rotated jumps select a unique attractive fixed point in our simulations; in practice, the relevant question is not exact microscopic invariance of the projected manifold, but whether off-manifold processes produce measurable steady-state deviations in accessible sizes and times, which they do not in the regime we consider.
@@ IV Dynamics: buildup of ODLRO and convergence to an $\eta$-paired NESS
The central claim of this work is that the rotated local dissipation stabilizes a nonequilibrium superconducting steady state starting from the fermionic vacuum
@@
in the precise sense of off-diagonal long-range order (ODLRO) in the $\eta$-pair channel. This choice of initial state is essential: $|\mathrm{vac}\rangle$ lies entirely in the holon–doublon doublet on every site and therefore provides a clean entry point for the dissipative pumping mechanism analyzed in Sec. II and the symmetry-based reduction of Sec. III. In particular, the protocol does not rely on fine-tuning a preformed paired state; rather, coherence is built up dynamically by the local jumps, which select a transverse $\eta$-pseudospin axis and actively phase-lock the system into an $\eta$-paired condensate.
@@
We diagnose pairing coherence using the local $\eta$-pair operator $\eta_{i}^{+}$ (and its Hermitian conjugate $\eta_{i}^{-}$). The local pairing amplitude,
@@
Figure 2: Real-time buildup of $\eta$-pair coherence under open boundary conditions (OBC) for $N=4$ at $t=\gamma$, $U=8/\gamma$, and the rotation angle $\theta=\pi/2$. Top panels: site-resolved one-point amplitude $|\Phi_{i}(\tau)|=|\langle\eta_{i}^{+}(t)\rangle|$ for (a1) a single driven site $\mathcal{J}=\{1\}$ and (a2) two driven sites $\mathcal{J}=\{1,4\}$; the dashed line indicates the projected benchmark $|\Phi_{i}|=1/2$. Bottom panels: real-space profile of the two-point correlator magnitude $|C(r,\tau)|=|\langle\eta_{i}^{+}(t)\eta_{i+r}^{-}(t)\rangle|$ for multiple separations $r$ (here $r=1,2,3$), shown for (b1) $\mathcal{J}=\{1\}$ and (b2) $\mathcal{J}=\{1,4\}$; the dashed line marks the predicted ODLRO plateau $|C|=1/4$. At intermediate times short-range correlations develop first, while at late times the correlator becomes nearly independent of $r$ and converges to an $r$-independent plateau $|C(r,t\!\to\!\infty)|\simeq 1/4$ (up to finite-size effects), providing direct evidence for ODLRO according to Eq. (31).
At the level of the leading projected generator (22), the steady state is the product dark state (24), i.e., a local $\eta$-pseudospin polarization along +y+y on every site. Equivalently, this NESS can be viewed as a coherent $\eta$-spin state obtained by rotating the trivial holon vacuum ($M=0$) into the +y+y direction:
@@
so the long-time state is not built by Hamiltonian pairing but rather selected as a Lindbladian attractor by the rotated local jumps. This immediately implies the projected benchmarks
$\left|\langle\eta_{i}^{+}\rangle_{\mathrm{ss}}\right|=\frac{1}{2},\qquad\left|\langle\eta_{i}^{+}\eta_{j}^{-}\rangle_{\mathrm{ss}}\right|=\frac{1}{4}\quad(i\neq j),$ (34)
i.e. an ODLRO plateau $C_{\infty}=1/4$ (up to finite-size corrections). Beyond the strictly projected picture, the full microscopic evolution admits perturbative excursions outside the projected manifold. These arise from (i) intra-holon–doublon mixing among different global-$\eta$ sectors induced by local $L_{j}$, and (ii) virtual leakage to singly occupied configurations mediated primarily by the hopping $H_{t}$. As discussed in Sec. III, such effects enter as higher-order corrections in the Schur-complement elimination of off-manifold components and predominantly renormalize transient relaxation pathways, while the leading attractive fixed point and the dominant decay scale $\Delta=\gamma/2$ [Eq. (26)] remain intact in the regime explored below.
We now verify these predictions by direct Lindblad time evolution and by computing $\langle O\rangle_{\tau}=\mathrm{Tr}[\rho(\tau)O]$. Figures 2(a1) and (a2) show the time evolution of the local $\eta$-pair amplitude $|\Phi_{i}(\tau)|\equiv|\langle\eta_{i}^{+}\rangle|$ and of a representative two-point correlator $|C(r,\tau)|$ for several choices of the driven set $\mathcal{J}$. In all cases the dynamics exhibits a rapid approach to the projected values, with $|\Phi_{i}(\tau\rightarrow\infty)|\rightarrow 1/2$ on a timescale governed by the dissipative gap. More importantly, Figures 2(b1) and (b2) display the full real-space profile $|C(r,\tau)|$ for multiple separations $r$, providing a direct visualization of the buildup of ODLRO. At intermediate times short-range correlations develop first, whereas at late times the correlator becomes nearly independent of $r$ and converges to an $r$-independent plateau, $|C(r,\tau\rightarrow\infty)|\simeq 1/4$ for all $1\leq r\leq N-1$ (up to finite-size effects). This long-distance saturation constitutes direct evidence for ODLRO according to Eq. (31).
@@
Before ending this part, we note that the resulting superconducting order is essentially a nonequilibrium phenomenon. The NESS does not emerge from thermalization under $H_{\mathrm{Hub}}$: within the degenerate $\eta$-multiplet $\mathcal{H}_{\Psi}$, the Hamiltonian itsself does not prefer a specific coherent orientation. Instead, the rotated local jumps act as a dissipative axis selector and continuously phase-lock the $\eta$ pseudospins, making the ODLRO plateau an attractor of the open-system dynamics. This picture aligns with symmetry considerations: since the local dissipation generally breaks the global $\eta^{2}$ [Eq. (15)] symmetry, the steady state cannot be viewed as an equilibrium (or generalized-Gibbs) ensemble dictated by the conserved charges of $H_{\mathrm{Hub}}$. Rather, ODLRO persists because it is actively stabilized by the driven dissipative dynamics, realizing a genuinely nonequilibrium $\eta$-paired steady state.
Figure 3: Finite-size scaling of global $\eta$-pair order parameters for dissipation applied only on the first site $\mathcal{J}=\{1\}$. (a) Global $\eta$-pair amplitude $|\Phi(\tau)|$. (b) $\eta$-pair structure factor $|S_{\eta}(\tau)|$. Curves are shown for different system sizes (as labeled). Both observables rapidly approach steady plateaus that are essentially insensitive to $N$, indicating that the observed macroscopic order is not a finite-size artifact. In particular, the saturation of $|S_{\eta}(\tau)|$ to a finite value, consistent with the real-space long-distance ODLRO limit, provides a complementary, bulk-sensitive confirmation of the dissipatively stabilized $\eta$-paired NESS. Parameters: 1D chain, $t=\gamma$, $U=4/\gamma$, and $\theta=\pi/2$.

### v2

# Single-site dissipation stabilizes a superconducting nonequilibrium steady state in a strongly correlated system

@@ Abstract
Can superconducting order be engineered as a robust attractor of open-system dynamics in strongly correlated systems? We demonstrate this possibility by proposing a minimal dissipation-engineering protocol for the particle-hole symmetric Hubbard model. By applying a rotated quantum jump operator, specifically a locally transformed $\eta$-pair lowering operator, on a single lattice site only, we show that the Lindblad evolution autonomously pumps the system from the vacuum into a nonequilibrium steady state (NESS) exhibiting macroscopic $\eta$-pair off-diagonal long-range order (ODLRO). Crucially, this local-to-global synchronization stands in stark contrast to schemes reliant on spatially extensive reservoirs: here, a single local dissipative seed suffices to establish long-range coherence across the entire interacting lattice system. We elucidate the underlying mechanism via three core features: local dark-state selection, the controlled elimination of off-manifold excursions induced by hopping, and a Liouvillian invariant-subspace structure that yields an attractive fixed point with a finite dissipative gap. Furthermore, we systematically classify the stability of this NESS with respect to static disorder, and identify a wide regime in which the superconducting attractor remains robust against Hamiltonian perturbations that preserve the effective subspace structure. We also pinpoint specific perturbations that directly cause dephasing of the $\eta$-pseudospin coherence and suppress ODLRO. These findings open up a disorder-tolerant pathway for stabilizing superconducting order as a non-thermal attractor through minimal local quantum-jump control.
@@ I Introduction
The paradigm of quantum state engineering has shifted fundamentally in recent years: rather than treating dissipation solely as a mechanism of decoherence, it is increasingly viewed as a resource for autonomous state preparation and stabilization[20, 16, 6, 11, 24, 19, 29, 23, 10]. By engineering the coupling between a system and its environment, one can design Liouvillian dynamics whose attractor (dark state) is a desired quantum many-body state[26, 27, 22, 31]. This approach has successfully demonstrated the stabilization of entangled states[18, 3, 25], topological phases[13, 2, 17, 15], and superfluid correlations in various atomic, molecular, and optical (AMO) and solid-state platforms[12, 14, 9].
However, a major open challenge lies in extending these protocols to strongly correlated systems where competing orders and constraints make the phase diagram fragile. Most existing schemes for stabilizing long-range order rely on spatially extensive dissipation—engineering reservoirs that couple to every site or bond of the lattice[12, 29]. While theoretically powerful, such bulk engineering is experimentally demanding and often incompatible with the natural local addressing capabilities of quantum devices. This raises a fundamental question of control and scalability: Is it possible to stabilize a macroscopic, strongly correlated quantum phase using only a minimal, strictly local dissipative seed?
In this work, we answer this question in the affirmative by demonstrating the dissipative stabilization of off-diagonal long-range order (ODLRO) in the Hubbard model using a single local dissipator. We focus on the paradigmatic setting of $\eta$-pairing superconductivity[30], an exact eigenstate phenomenon in the Hubbard model that implies distance-independent pairing correlations protected by an SU(2) pseudospin symmetry. While $\eta$-pairing states are typically high-energy excited states in equilibrium and are thus thermodynamically unstable[28, 21],they represent ideal candidates for non-equilibrium steady states (NESS) in the presence of appropriate dynamical symmetries [5, 8].
We introduce a protocol based on a rotated quantum jump operator acting on as few as a single lattice site. Our results reveal a striking local-to-global mechanism: this local dissipative seed effectively pumps the entire interacting many-body system into a specific holon–doublon manifold, triggering a self-organization process that establishes global superconducting coherence. Crucially, this mechanism bypasses the need for bulk reservoir engineering. Instead, the jump operator’s dark-state structure combines with coherent Hamiltonian dynamics to propagate the locally injected pseudospin coherence across the lattice.
@@
The rest of the paper is organized as follows. In Sec. II, we provide a heuristic derivation that distills the main idea of the rotated-jump scheme and its local dark-state structure. In Sec. III, we present the full protocol for the Hubbard chain and analyze the underlying mechanism from the viewpoint of symmetry, with emphasis on the Liouvillian structure and invariant subspaces. Section IV is devoted to the dynamical diagnostics of superconducting behavior: we numerically track the time evolution of two pairing correlators and demonstrate the emergence of $\eta$-pair ODLRO in the NESS. In Sec. V, we investigate disorder effects in a systematic manner and delineate robust versus destructive perturbations. Finally, Sec. VI summarizes our findings and outlines future directions. Technical derivations and additional details are provided in the Appendices.
Figure 1: Schematic of the non-equilibrium phase-locking protocol that stabilizes an $\eta$-paired superconducting NESS. Starting from a strongly correlated Hubbard system (with optional disorder in the interaction $U_{i}$, onsite potential $\mu_{i}$, and/or bond hopping $t_{ij}$), we apply local rotated dissipation on selected sites with jump operators $L_{j}=\eta_{j}^{-}e^{i\frac{\pi}{2}\eta_{j}^{x}}$. Unlike the unrotated channel $\tilde{L}_{j}=\eta_{j}^{-}$, which simply annihilates pairs and leaves the vacuum as the local dark state, the rotation tilts the dissipative measurement axis so that the locally protected state is the +y+y $\eta$-pseudospin eigenstate, $|\eta_{j}^{y}=+1/2\rangle\propto e^{-i\frac{\pi}{2}\eta_{j}^{x}}|0\rangle_{j}$, i.e., a coherent holon–doublon superposition. Each quantum jump then acts like a small compass correction: it damps misaligned components while repumping transverse $\eta$ coherence. In the presence of strong correlations, this local axis selection propagates through the many-body dynamics, effectively phase-locking the $\eta$ pseudospins across the lattice. At long times, the system relaxes to a unique NESS with $\eta$ pseudospins aligned along the +y+y direction, yielding a finite $\eta$-pair amplitude and ODLRO.
@@ II Single-qubit heuristic: rotated dissipation pumps coherence from the vacuum
and set $H=0$ throughout this section. The goal of this toy example is to make one mechanism completely transparent, in a form that will reappear in the Hubbard chain: a jump operator does not merely relax the system, it selects an attractive fixed point (a dark state or dark manifold). Consequently, rotating the jump operator rotates the dissipatively selected axis and thus the steady state itself. In this way, dissipation acts not as a destroyer of phase information, but rather as a simple coherence pump.
@@
Equivalently, the steady-state Bloch vector points along $\hat{y}$, with $\langle s^{x}\rangle=0$ and $\langle s^{y}\rangle=1/2$. Under the mapping (3), Eq. (6) is precisely the single-site analogue of $\langle\eta_{i}^{+}\rangle\neq 0$: the rotated dissipation does not create coherence out of nothing; rather, it reshapes the dark-state structure so that the unique attractor contains the desired coherence between $|0\rangle$ and $|\uparrow\downarrow\rangle$.
This local picture foreshadows the many-body mechanism exploited below. In the Hubbard chain, implementing the analogous rotated jump in the $\eta$-spin sector selects locally coherent holon–doublon superpositions as building blocks. The remaining task is to understand how interactions and hopping propagate and phase-lock this locally seeded $\eta$-pseudospin orientation across the lattice, resulting in an $\eta$-paired NESS with off-diagonal long-range order.
@@ III.4 Leading projected Liouvillian, NESS, and the dissipative gap
To address potential objections, we offer three key clarifications. (i) Effective gap and timescales: Eq. (26) describes a property of the projected effective Liouvillian–specifically, that the leading-order reduced generator exhibits a finite spectral separation scale set by the local pumping strength. In the full microscopic model, additional processes (e.g., hopping-assisted excursions and local sector mixing) renormalize relaxation pathways and hinder the clean extraction of a single-exponential rate. Rather than overinterpreting fitted microscopic gap values, we adopt $\Delta\sim\gamma/2$ as an organizing scale that accounts for the rapid convergence and suppressed deviations within the strong-coupling/controlled regime where effective elimination holds. (ii)Finite-size scaling: To distinguish genuine long-range order from finite-size saturation, we track the $\eta$-pair structure factor $S_{\eta}(\tau)$ as well as the global amplitude $|\Phi(\tau)|$ for progressively larger system sizes, with results presented in the next subsection. The weak size dependence exhibited by these steady-state plateaus confirms an $\mathcal{O}(1)$ order parameter in the thermodynamic limit. (iii) Uniqueness within the studied sector: Starting from the vacuum (holon–doublon) sector, the rotated jumps yield a unique attractive fixed point in our simulations. In practice, the key question is not whether the projected manifold is exactly microscopically invariant, but whether off-manifold processes induce measurable steady-state deviations in accessible system sizes and times, and in the regime we consider, such deviations are absent.
@@ IV Dynamics: buildup of ODLRO and convergence to an $\eta$-paired NESS
The key claim of this work is that rotated local dissipation stabilizes a non-equilibrium superconducting steady state emerging from the fermionic vacuum,
@@
specifically as defined by ODLRO in the $\eta$-pair channel. This choice of initial state is essential: $|\mathrm{vac}\rangle$ lies entirely in the holon–doublon doublet on every site and therefore provides a clean entry point for the dissipative pumping mechanism analyzed in Sec. II and the symmetry-based reduction of Sec. III. In particular, the protocol does not rely on fine-tuning a preformed paired state; rather, coherence is built up dynamically by the local jumps, which select a transverse $\eta$-pseudospin axis and actively phase-lock the system into an $\eta$-paired condensate.
@@
We examine pairing coherence using the local $\eta$-pair operator $\eta_{i}^{+}$ (and its Hermitian conjugate $\eta_{i}^{-}$). The local pairing amplitude,
@@
Figure 2: Real-time buildup of $\eta$-pair coherence under open boundary conditions (OBC) for $N=4$ at $t=\gamma$, $U=8\gamma$, and the rotation angle $\theta=\pi/2$. Top panels: site-resolved one-point amplitude $|\Phi_{i}(\tau)|=|\langle\eta_{i}^{+}(t)\rangle|$ for (a1) a single driven site $\mathcal{J}=\{1\}$ and (a2) two driven sites $\mathcal{J}=\{1,4\}$; the dashed line indicates the projected benchmark $|\Phi_{i}|=1/2$. Bottom panels: real-space profile of the two-point correlator magnitude $|C(r,\tau)|=|\langle\eta_{i}^{+}(t)\eta_{i+r}^{-}(t)\rangle|$ for multiple separations $r$ (here $r=1,2,3$), shown for (b1) $\mathcal{J}=\{1\}$ and (b2) $\mathcal{J}=\{1,4\}$; the dashed line marks the predicted ODLRO plateau $|C|=1/4$. At intermediate times, short-range correlations develop first, while at late times the correlator becomes nearly independent of $r$ and converges to an $r$-independent plateau $|C(r,t\!\to\!\infty)|\simeq 1/4$ (up to finite-size effects), providing direct evidence for ODLRO according to Eq.  (31).
At the level of the leading projected generator (22), the NESS can be viewed equivalently as a coherent $\eta$-spin state obtained by rotating the trivial holon vacuum ($M=0$) into the +y+y direction
@@
so the long-time state is not built by Hamiltonian pairing but rather selected as a Lindbladian attractor by the rotated local jumps. This directly gives the projected benchmark, namely an ODLRO plateau of $C_{\infty}=1/4$ (up to finite-size corrections). Beyond the strictly projected picture, the full microscopic evolution admits perturbative excursions outside the projected manifold. These arise from (i) intra-holon–doublon mixing among different global-$\eta$ sectors induced by local $L_{j}$, and (ii) virtual leakage to singly occupied configurations mediated primarily by the hopping $H_{t}$. As discussed in Sec. III, such effects enter as higher-order corrections in the Schur-complement elimination of off-manifold components and predominantly renormalize transient relaxation pathways, while the leading attractive fixed point and the dominant decay scale $\Delta=\gamma/2$ [Eq. (26)] remain intact in the regime explored below.
We now verify these predictions via direct Lindblad time evolution and explicit computation of $\langle O\rangle_{\tau}=\mathrm{Tr}[\rho(\tau)O]$. Figures 2(a1) and (a2) show the time evolution of both the local $\eta$-pair amplitude $|\Phi_{i}(\tau)|\equiv|\langle\eta_{i}^{+}\rangle|$ and a representative two-point correlator $|C(r,\tau)|$ for multiple choices of the driven set $\mathcal{J}$. In all cases, the dynamics exhibits a rapid approach to the projected values, with $|\Phi_{i}(\tau\rightarrow\infty)|\rightarrow 1/2$ on a timescale governed by the dissipative gap. More importantly, Figures 2(b1) and (b2) display the full real-space profile $|C(r,\tau)|$ for multiple separations $r$, providing a direct visualization of the buildup of ODLRO. At intermediate times short-range correlations develop first, whereas at late times the correlator becomes nearly independent of $r$ and converges to an $r$-independent plateau, $|C(r,\tau\rightarrow\infty)|\simeq 1/4$ for all $1\leq r\leq N-1$ (up to finite-size effects). This long-distance saturation constitutes direct evidence for ODLRO according to Eq. (31).
@@
Before ending this part, we note that the resulting superconducting order is essentially a non-equilibrium phenomenon. The NESS does not emerge from thermalization under $H_{\mathrm{Hub}}$: within the degenerate $\eta$-multiplet $\mathcal{H}_{\Psi}$, the Hamiltonian itsself does not prefer a specific coherent orientation. Instead, the rotated local jumps act as a dissipative axis selector and continuously phase-lock the $\eta$ pseudospins, making the ODLRO plateau an attractor of the open-system dynamics. This picture aligns with symmetry considerations: since the local dissipation generally breaks the global $\eta^{2}$ [Eq. (15)] symmetry, the steady state cannot be viewed as an equilibrium (or generalized-Gibbs) ensemble dictated solely by the conserved charges of $H_{\mathrm{Hub}}$. Rather, ODLRO persists because it is actively stabilized by the driven dissipative dynamics, realizing a genuinely non-equilibrium $\eta$-paired steady state.
Figure 3: Finite-size scaling of global $\eta$-pair order parameters for dissipation applied only on the first site $\mathcal{J}=\{1\}$. (a) Global $\eta$-pair amplitude $|\Phi(\tau)|$. (b) $\eta$-pair structure factor $|S_{\eta}(\tau)|$. Curves are shown for different system sizes (as labeled). Both observables rapidly approach steady plateaus that are essentially insensitive to $N$, indicating that the observed macroscopic order is not a finite-size artifact. In particular, the saturation of $|S_{\eta}(\tau)|$ to a finite value, consistent with the real-space long-distance ODLRO limit, provides a complementary, bulk-sensitive confirmation of the dissipatively stabilized $\eta$-paired NESS. Parameters: 1D chain, $t=\gamma$, $U=4\gamma$, and $\theta=\pi/2$.